OpenStudy (anonymous):

What value of k solves the equation? k^-4 = 1/256

OpenStudy (slaaibak):

Hint: 256 = 2^8 = (-2)^8

OpenStudy (anonymous):

whats the answer

OpenStudy (slaaibak):

What did you get?

OpenStudy (anonymous):

ok i got -64

jhonyy9 (jhonyy9):

so if you check it right,sure ,correct and will make the calcule will get k=4

OpenStudy (slaaibak):

Nope. \[{1 \over k^4 } ={ 1 \over 256} \rightarrow k^4 = 256\]

OpenStudy (slaaibak):

^jhonyy9, don't forget about k=-4

jhonyy9 (jhonyy9):

check it ,,slaaibak" !!!

OpenStudy (anonymous):

is it - or +

OpenStudy (slaaibak):

both.

jhonyy9 (jhonyy9):

yes +/- 4 will be right